OpenStudy (anonymous):

find extrema of f(x) = x + 2cosx on the interval [-pi, pi] I got x = pi/6, and x=5pi/6. One is a min, and one is a max, but the book also got x= -2. Can someone explain how they got this?

OpenStudy (anonymous):

lets check it\[f'(x)=1-2\sin x=0\]gives\[\sin x=\frac{1}{2}\]x=-2 cant be one of extremas

OpenStudy (anonymous):

ithink what u got is right :)

OpenStudy (anonymous):

@mukushla This is a screenshot of the answer from the textbook. I'm not sure where the two came from, but it is somehow a part of the answer -_-;; Thanks for replying though XD

OpenStudy (anonymous):

yw :)